resultant of two equal forces is double of either of force the angle b/w them is
ans.zero degree

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
R^2 = p^2 + q^2 + 2 pq cos@
 
4p^2 = p^2 + p^2 + 2p^2 cos @
 
2p^2 cos@ = 2p^2
 
hence cos @= 1
 
so
 
@ = 0 degree
 
Regards
Arun